OpenStudy (anonymous):

you and your friend spent a total of $15 for lunch. your friend's lunch cost $3 more than your cost. how much did you spend for lunch?

OpenStudy (noseboy908):

6 dollars

OpenStudy (anonymous):

$6

OpenStudy (p0sitr0n):

your lunch = x your friend's = x+3 total = 15 x + x + 3 = 15 2x +3 = 15 2x = 12 x=6$

OpenStudy (anonymous):

showoff^ lol